# On the correlation between entanglement and the negative sign problem

### v1

@@ I Introduction
At the first glance, it seems that entanglement has nothing to do with the average sign from their definitions. However, previous studies suggest that there indeed exists correlation between them. It is known that the anti-ferromagnetic Heisenberg model on bipartite lattices (e.g., one-dimensional, square, honeycomb lattices) is sign problem free. At the same time, these systems are easier to study with tensor network state related methods than the Heisenberg model on frustrated lattices such as triangular [15] and kagome [16, 17, 18] lattices, which encounter negative sign problem in QMC. In [19], it is found that a typical random positive wave-function (which is in principle sign problem free) can’t support volume law entanglement. The diffusion Monte Carlo (DMC) method can accurately simulate one-dimensional (1D) fermion system because the node structure of 1D fermion ground state is solely determined by the minus sign from the anti-commutation relation of fermion [20]. Interestingly, it is well known that Density Matrix Renormalization Group (DMRG) [8] can easily handle 1D quantum system because the entanglement entropy in the ground state of 1D system scales at most logarithmically with the size of the system. However, both DMC and DMRG encounter difficulty in the study of 2D systems. On the one hand, the sign structure of 2D quantum system is far more complicated than the 1D case which makes it is almost impossible to determine the exact sign structure of general system [20, 21, 22]. This can be viewed as a manifestation of the negative sign problem. On the other hand, area law [23] tells that the entanglement entropy of the ground state of 2D systems scales at least as the linear dimension of the systems, making the study of 2D quantum system with DMRG difficult. We notice that there also exist 2D tensor network state methods [10, 11, 12, 13], but the heavy cost makes the entanglement they can handle limited. In [24], it is found that sign-free model can’t have a stable Fermi surface, while it is known that the Fermi system with non-trivial Fermi surface usually causes a logarithm correlation to the area law of entanglement [25, 26]. This means that system with negative sign problem in QMC usually has large entanglement in the ground state. In [27], it is also shown that sign problem is quantitatively linked to quantum critical behavior, and it is known that the entanglement usually increases when approaching quantum critical point [28, 29]. Recent study also shows that the positive definiteness of the elements of tensor network makes the contraction of it easier [30]. We also notice that the average sign is a contribution term in the second order Renyi entropy defined with the replica trick [31, 32, 33].
@@
In this work, we take the Hubbard model with size $4\times 16$ and $U/t=8$ as an example to study the correlation between entanglement and average sign. We carefully study the doping evolution of the entanglement in the ground state with DMRG and the average sign in QMC at low temperature. We find that they are indeed correlated. The entanglement (average sign) shows a peak (dip) around $20\%$ doping, which means that it is the most difficult region for both QMC and tensor network state related methods. We also notice that the vicinity of $20\%$ doping is the most intriguing region in both Hubbard model and cuprates where competing states with close energies intertwine with each other [7, 34, 35, 36]. These results indicate that systems which are difficult to simulate with many-body methods probably host interesting physics.
@@ II.3 Average sign
and Hubbard Stratonovich transformation in the charge decomposition form as
$e^{-\Delta\tau Un_{i\uparrow}n_{i\downarrow}}=e^{-\Delta\tau U\left(n_{i\uparrow}+n_{i\downarrow}-1\right)/2}\sum_{s_{i}=\pm 1}\frac{1}{2}e^{\gamma s_{i}\left(n_{i\uparrow}+n_{i\downarrow}-1\right)}$ (5)
where $\cosh(\gamma)=\exp(-\Delta\tau U/2)$ and $s_{i}$ is the introduced auxiliary field. We denote ${V_{s}}=-\gamma\sum\limits_{i}{{s_{i}}({n_{i\uparrow}}+{n_{i\downarrow}})}+\Delta\tau U\sum\limits_{i}{({n_{i\uparrow}}+{n_{i\downarrow}})/2}$.
@@
$Z={\rm{Tr}}({e^{-\beta H}})=\sum\limits_{\{s\}}{{Z_{s}}}=\sum\limits_{\{s\}}{{C_{\{s\}}}\det(1+{B_{s}})}$ (6)
where ${B_{s}}=\prod\limits_{m=1}^{\beta/\Delta\tau}{{e^{-K\Delta\tau/2}}{e^{-{V_{s(m)}}}}{e^{-K\Delta\tau/2}}}$ and ${{C_{\{s\}}}}$ is the constant coefficient.
@@
where $\left\langle O_{s}\right\rangle$ is the expectation value of the operator ${\hat{O}}$ for a single auxiliary field configuration and $\left\langle{{\mathop{\rm sgn}}}\right\rangle$ is the average sign:
@@
[…]
@@ III Results
We study the Hubbard model on a $4\times 16$ cylinder with $U/t=8$. We vary the filling factor and study a range of doping from zero doping (half-filling) to doping around $h=0.5$. We carefully examine the doping evolution of both the average sign in AFQMC at low temperature and the entanglement and truncation error in DMRG calculations of the ground state.
@@
Figure 1: The evolution of (a) average sign in AFQMC, (b) entanglement entropy and (c) truncation error in DMRG with doping. In (a), average sign for $\beta=2.5$ and $3$ are shown. In (c), the truncation errors in DMRG of different bond dimensions are shown. For the half-filling and the $h=0.0625$ cases, small bond dimension already gives very small truncation error. We can find a correlation between average sign and entanglement entropy or truncation error. By doping away from the half-filling case, the average sign first decreases, reaching its minimum value at about $20\%$ doping, then starts to increase with the further increase of doping. The entanglement entropy shown in (b) are the values from a linear extrapolation with truncation error. The entanglement entropy and truncation error have opposite behavior as the average sign. Notice that region for small average sign and large entanglement/truncation error are difficult for both AFQMC and DMRG.
@@
We show the evolution of average sign in AFQMC, entanglement entropy and truncation error in DMRG with doping in Fig. 1. The truncation error in DMRG can be used to judge the accuracy of the calculation. In AFQMC calculation, we study two temperatures with $\beta=2.5$ and $3$. Similar results can be found in many studies in the literature [41, 14, 46]. We can clearly see that the average sign decreases by lowing the temperature as expected. Af half-filling, the Hubbard model is sign problem free with $\langle\text{sign}\rangle=1$ as shown in Fig. 1. We find that the average sign first decreases when the system is doped away from half filling, reaching its minimum value around $20\%$ doping, and then starts to increase with doping. The entanglement entropy in Fig. 1 (b) are obtained from a linear extrapolation of the results with truncation error, whose values are plot in Fig. 1 (c) for different bond dimensions. For the half-filling and the $h=0.0625$ cases, small bond dimension already gives very small truncation error. The entanglement entropy and truncation error have opposite behavior as the average sign. Half filling system has the smallest entanglement entropy and smallest truncation error with fixed bond dimension. With the increase of doping, the entanglement entropy and truncation error first increase and reach the maximum value also around $20\%$ doping, and then start to decrease. The results show that the region around $20\%$ doping is the most difficult region for both DMRG and AFQMC methods. Experimentally, $20\%$ doping is also the most intriguing region where different states close in energy intertwine with each other [7]. These results indicate that systems which are difficult to simulate with many-body methods probably host interesting physics.
@@
Our discussion so far has focused on the “exact” quantum Monte Carlo method. We notice that there also exist quantum Monte Carlo methods [56, 57] where approximations are adopted to eliminate the negative sign problem. The essence of these methods is a trade-off between the negative sign problem and systematic error. Results has shown that the systematic error is mild in many cases [58, 2]. It will be interesting to know whether there is a fundamental correlation between systematic error and entanglement since there is freedom in choosing the approximation to eliminate the negative sign problem.

### v2

@@ I Introduction
At the first glance, it seems that entanglement has nothing to do with the average sign from their definitions. However, previous studies suggest that there indeed exists correlation between them. It is known that the anti-ferromagnetic Heisenberg model on bipartite lattices (e.g., one-dimensional, square, honeycomb lattices) is sign problem free. At the same time, these systems are easier to study with tensor network state related methods than the Heisenberg model on frustrated lattices such as triangular [49] and kagome [11, 55, 28] lattices, which encounter negative sign problem in QMC. In [16], it is found that a typical random positive wave-function (which is in principle sign problem free) can’t support volume law entanglement. The diffusion Monte Carlo (DMC) method can accurately simulate one-dimensional (1D) fermion system because the node structure of 1D fermion ground state is solely determined by the minus sign from the anti-commutation relation of fermion [6]. Interestingly, it is well known that Density Matrix Renormalization Group (DMRG) [50] can easily handle 1D quantum system because the entanglement entropy in the ground state of 1D system scales at most logarithmically with the size of the system. However, both DMC and DMRG encounter difficulty in the study of 2D systems. On the one hand, the sign structure of 2D quantum system is far more complicated than the 1D case, making it almost impossible to determine the exact sign structure of general system [6, 24, 23]. This can be viewed as a manifestation of the negative sign problem. On the other hand, area law [10] tells that the entanglement entropy of the ground state of 2D systems scales at least as the linear dimension of the systems, making the study of 2D quantum system with DMRG difficult. We notice that there also exist 2D tensor network state methods [34, 8, 3, 53], but the heavy cost makes the entanglement they can handle limited. In [15], it is found that sign problem-free model can’t have a stable Fermi surface, while it is known that the Fermi system with non-trivial Fermi surface usually causes a logarithm correlation to the area law of entanglement [42, 51]. This means that system with negative sign problem in QMC usually has large entanglement in the ground state. In [31, 30], it is also shown that sign problem is linked to quantum critical behavior, and it is known that the entanglement usually increases when approaching quantum critical point [45, 59]. Recent study also shows that the positive definiteness of the elements of tensor network makes the contraction of it easier [7]. We also notice that the average sign is a contribution term in the second order Renyi entropy defined with the replica trick [18, 4, 5].
@@
In this work, we take the Hubbard model with size $4\times 16$ and $U/t=8$ as an example to study the correlation between entanglement and average sign. Our focus is the ground state or low temperature simulations. We carefully study the doping evolution of the entanglement in the ground state with DMRG and the average sign in QMC at low temperature. We find that they are indeed correlated. The entanglement (average sign) shows a peak (dip) around $20\%$ doping, which means that it is the most difficult region for both QMC and tensor network state related methods. We also notice that the vicinity of $20\%$ doping is the most intriguing region in both Hubbard model and cuprates where competing states with close energies intertwine with each other [13, 37, 60, 54]. These results indicate that systems which are difficult to simulate with many-body methods probably host interesting physics.
@@ II.3 Average sign
and Hubbard Stratonovich transformation in the spin decomposition form as
$e^{-\Delta\tau Un_{i\uparrow}n_{i\downarrow}}=e^{-\Delta\tau U(n_{i\uparrow}+n_{i\downarrow})/2}\sum_{s_{i}=\pm 1}\frac{1}{2}e^{\gamma s_{i}(n_{i\uparrow}-n_{i\downarrow})}$ (5)
where $\cosh(\gamma)=\exp(\Delta\tau U/2)$ and $s_{i}$ is the introduced auxiliary field. We denote ${V_{s}}=-\gamma\sum\limits_{i}{{s_{i}}({n_{i\uparrow}}-{n_{i\downarrow}})}+\Delta\tau U\sum\limits_{i}{({n_{i\uparrow}}+{n_{i\downarrow}})/2}$.
@@
$Z={\rm{Tr}}({e^{-\beta H}})=\sum\limits_{\{s\}}{{Z_{s}}}=\sum\limits_{\{s\}}{\det(1+{B_{s}})}$ (6)
where ${B_{s}}=\prod\limits_{m=1}^{\beta/\Delta\tau}{{e^{-K\Delta\tau/2}}{e^{-{V_{s(m)}}}}{e^{-K\Delta\tau/2}}}$.
@@
where $\left\langle O_{s}\right\rangle$ is the expectation value of the operator $^{\hat{O}}$ for a single auxiliary field configuration and $\left\langle{{\mathop{\rm sgn}}}\right\rangle$ is the average sign:
@@
For general system with negative sign problem, the average sign usually scale as $\langle\text{sign}\rangle\sim\exp(-c\beta N)$ where $\beta$ is the inverse of temperature, $N$ is the lattice size and $c$ is a constant [29]. From Eq. (8), we can find that the observable $\left\langle O\right\rangle$ has huge fluctuation at low temperature in large systems. So the value of $\langle\text{sign}\rangle$ can be utilized as a measure of the difficulty in QMC calculation. Although it is known that there are cases where the negative sign problem is absent [20, 52, 47, 46, 27], general systems usually suffer from the negative sign problem.
@@ III Results
We study the Hubbard model on a $4\times 16$ cylinder (with periodic (open) boundary conditions along the short (long) direction) with $U/t=8$. We use AFQMC and DMRG to calculate the exact same systems. We vary the filling factor and study a range of doping from zero doping (half-filling) to doping around $h=0.5$. We carefully examine the doping evolution of both the average sign in AFQMC at low temperature and the entanglement and truncation error in DMRG calculations of the ground state.
@@
Figure 1: The evolution of (a) average sign in AFQMC, (b) entanglement entropy and (c) truncation error in DMRG with doping. In (a), average sign for $\beta=2.5$, $3$ and $4$ are shown. In (c), the truncation errors in DMRG of different bond dimensions are shown. For the half-filling and the $h=0.0625$ cases, small bond dimension already gives very small truncation error. We can find a correlation between average sign and entanglement entropy or truncation error. By doping away from the half-filling case, the average sign first decreases, reaching its minimum value at about $20\%$ doping, then starts to increase with the further increase of doping. The entanglement entropy shown in (b) are the values from a linear extrapolation with truncation error. The entanglement entropy and truncation error have opposite behavior as the average sign. Notice that region for small average sign and large entanglement/truncation error are difficult for both AFQMC and DMRG.
@@
We show the evolution of average sign in AFQMC, entanglement entropy and truncation error in DMRG with doping in Fig. 1. The truncation error in DMRG can be used to judge the accuracy of the calculation. In AFQMC calculation, we study two temperatures with $\beta=2.5$ ,33 and $4$. Similar results can be found in many studies in the literature [20, 29, 48]. We can clearly see that the average sign decreases by lowing the temperature as expected. Af half-filling, the Hubbard model is sign problem free with $\langle\text{sign}\rangle=1$ as shown in Fig. 1. We find that the average sign first decreases when the system is doped away from half filling, reaching its minimum value around $20\%$ doping, and then starts to increase with doping. The entanglement entropy in Fig. 1 (b) are obtained from a linear extrapolation of the results with truncation error, whose values are plot in Fig. 1 (c) for different bond dimensions. For the half-filling and the $h=0.0625$ cases, small bond dimension already gives very small truncation error. The entanglement entropy and truncation error have opposite behavior as the average sign. Half filling system has the smallest entanglement entropy and smallest truncation error with fixed bond dimension. With the increase of doping, the entanglement entropy and truncation error first increase and reach the maximum value also around $20\%$ doping, and then start to decrease. The results show that the region around $20\%$ doping is the most difficult region for both DMRG and AFQMC methods. Experimentally, $20\%$ doping is also the most intriguing region where different states close in energy intertwine with each other [13]. These results indicate that systems which are difficult to simulate with many-body methods probably host interesting physics.
@@
Sign problem usually indicates the cancellation of positive and negative paths in the sampled space. The superposition of extensive positive and negative paths means massive destructive interference which usually results in highly entangled states. Physically, the competition between different candidate states in the low-energy manifold—whether due to frustration or the interplay between kinetic and potential energies—typically generates strong entanglement and a severe sign problem, leading to exotic quantum phenomena that are challenging to simulate classically. In this sense, average sign and entanglement entropy is naturally connected in challenging systems and can be used as an indicator to exotic physics [31, 30].
@@
Our discussion so far has focused on the ”exact” quantum Monte Carlo method. We notice that there also exist quantum Monte Carlo methods [56, 12] where approximations are adopted to eliminate the negative sign problem. The essence of these methods is a trade-off between the negative sign problem and systematic error. Results has shown that the systematic error is mild in many cases [39, 25]. It will be interesting to know whether there is a fundamental correlation between systematic error and entanglement since there is freedom in choosing the approximation to eliminate the negative sign problem.
@@ Acknowledgements.
Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.